# VLA-Pruner: Temporal-Aware Dual-Level Visual Token Pruning for Efficient Vision-Language-Action Inference

### v2

@@ Abstract
Vision-Language-Action (VLA) models have shown great promise for embodied AI, yet the heavy computational cost of processing continuous visual streams severely limits their real-time deployment. Token pruning—keeping salient visual tokens and dropping redundant ones—has emerged as an effective approach for accelerating Vision-Language Models (VLMs), offering a solution for efficient VLA. However, these VLM-specific token pruning methods select tokens based solely on semantic salience metrics (e.g., prefill attention), while overlooking the VLA’s intrinsic dual-system nature of high-level semantic understanding and low-level action execution. Consequently, these methods bias token retention toward semantic cues, discard critical information for action generation, and significantly degrade VLA performance. To bridge this gap, we propose VLA-Pruner, a versatile plug-and-play VLA-specific token prune method that aligns with the dual-system nature of VLA models and exploits the temporal continuity in robot manipulation. Specifically, VLA-Pruner adopts a dual-level importance criterion for visual token retention: vision-language prefill attention for semantic-level relevance and action decode attention, estimated via temporal smoothing, for action-level importance. Based on this criterion, VLA-Pruner proposes a novel dual-level token selection strategy that adaptively preserves a compact, informative set of visual tokens for both semantic understanding and action execution under given compute budget. Experiments show that VLA-Pruner achieves state-of-the-art performance across multiple VLA architectures and diverse robotic tasks.
@@ 1 Introduction
Recent advances in robot learning have led to remarkable progress in developing generalist policies capable of acting across diverse tasks and environments [35, 13, 62, 7, 63, 22, 4, 25, 9]. Vision–Language–Action (VLA) models [7, 63, 22, 23, 4, 25, 30], which benefit from the multimodal understanding and transfer ability of Vision–Language Models (VLMs) [33, 3], have emerged as a promising pathway toward general-purpose robot policy. Leveraging large-scale real-world robotic datasets [40], VLA models can perceive visual scenes, understand natural language instructions, and directly execute low-level actions to complete complex tasks across diverse scenarios. However, in real-time deployment, VLA models must process continuous visual streams within dynamic environments and thus struggle with heavy computational overhead. To mitigate this, existing works have explored acceleration techniques such as model lightweighting [51], quantization [41], and early-exit [59] for VLA. Despite effectiveness to some extent, these methods require either architectural modifications or retraining, therefore hindered by poor generalizability.
@@
Previous research [8, 47, 2] has shown that visual representations (i.e., visual tokens) in Vision-Language Models (VLMs) are highly redundant. Consequently, visual token pruning, which retains the most salient tokens while pruning the rest, has become a general and effective approach to improve VLM inference efficiency [2, 8, 60, 47, 57, 53, 64] and shows promise for general VLA acceleration. These methods typically define importance criteria for tokens, such as attention scores [8, 60, 64], gradient information [37, 36] or feature diversity [2] to quantify the significance of visual tokens, and less important tokens are pruned at inference time. By reducing visual redundancy, they yield substantial compute savings with minimal accuracy loss on standard VLM tasks such as image captioning [1, 44], VQA [20, 34], and video understanding [54, 28].
@@
Despite recent progress, existing token pruning methods typically rely solely on semantic salience metrics measured during context prefilling—e.g., prefill attention distributions in FastV [8] or text-to-vision cross-attention scores in SparseVLM [60]—to rank visual tokens and prune the lowest-scoring ones. While such semantic-only criteria can be effective for VLMs, they are misaligned with the intrinsic dual-system nature of VLA models [18, 19, 48, 11]. Specifically, VLA inference couples high-level semantic understanding, task planning with low-level action execution, which entails distinct visual requirements between its vision-language prefill and action decode stages; correspondingly, we observe clearly different attention patterns across these two stages (see Sec. 3.2). Consequently, current methods tend to bias token retention towards semantic cues, thereby discarding critical details for action generation and significantly degrading VLA performance, especially at high pruning ratios (see Fig. 1), as also observed in prior work [55]. Thus, existing VLM-specific token pruning strategies are fundamentally ill-suited for VLA models, highlighting the need for a VLA-specific method.
@@
To bridge this gap, we propose VLA-Pruner, a versatile plug-and-play VLA token pruning approach that aligns with the dual-system nature of VLA models and exploits temporal continuity in robot manipulation. VLA-Pruner offers two principal insights for VLA token pruning. First, VLA-Pruner adopts a dual-level importance criterion for visual token retention: vision-language prefill attention scores to quantify semantic-level relevance (following prior works [8, 57, 37, 60]), and action decode attention scores to quantify action-level importance. A critical challenge to achieve this is that decode attention is unavailable at prefill time. Fortunately, we observe that temporal continuity of robot manipulation [55] manifests in VLA action attention patterns evolving: top-attended visual patches overlap substantially across consecutive timesteps (see Sec. 3.2). Leveraging this continuity, we practically estimate the current action-to-vision attention from recent decode attentions via statistical temporal smoothing techniques (e.g., Exponential Moving Average [15, 14]). Second, to ensure VLA performance, our novel dual-level token selection strategy avoids fragile attention score fusion that incurs redundancy and bias. Instead, drawing on the minimal-redundancy-maximal-relevance (mRMR) principle [42], we adopt a patch-wise scheme: we first take the union of patches salient at either semantic or action level to maximize relevance, and then filter them by reducing redundancy, yielding a compact yet informative visual token set that meets the given compute budget. This innovative design reduces computational overhead and adaptively preserves essential information for both semantic understanding and action execution. VLA-Pruner is training-free, serving as a plug-and-play acceleration module for most state-of-the-art VLA architectures [22, 24, 4]. We evaluate VLA-Pruner on robotic manipulation across two simulated environments (LIBERO [32] and SIMPLER [27]) and three VLA models (OpenVLA [22], OpenVLA-OFT [24], and [4]). VLA-Pruner delivers substantial efficiency gains (up to 1.8 speedup) with minor drops in task success rate. Even at 87.5% prune ratio, VLA-Pruner can maintain reasonable performance. Moreover, thanks to precise redundancy removal, VLA-Pruner can improve model performance at the 50% prune ratio. Finally, we demonstrate its real-world applicability by deploying it on a 6-DoF xArm6 robot, achieving practical speedups in real-time scenarios.
2 Related Work
2.1 Vision-Language-Action Models (VLA)
@@
Large-scale vision–language models (VLMs) have propelled multimodal learning by incorporating visual perception and language reasoning [26]. Building on these advances, Vision–Language–Action (VLA) models [22, 4, 24, 7, 63] introduce an action modality, enabling end-to-end motor control. These models typically adopt large VLM backbones [49] and are fine-tuned on robotic data [40]. For action generation, implementations range from discretizing actions into language-like tokens (e.g., OpenVLA [22]) to attaching diffusion-policy heads (e.g., [4]), typically relying on action-to-vision cross-attention for fine-grained interaction. Despite their effectiveness on robot manipulation tasks such as object retrieval and assembly [32, 27], VLA models remain computationally intensive, hindering real-time deployment on resource-constrained platforms.
2.2 Visual Token Pruning for VLMs
@@
Visual token pruning [8, 60, 2, 47, 57] is a widely adopted strategy to reduce visual redundancy in VLMs. One line of methods defines semantic importance criteria to rank and prune visual tokens, e.g., FastV [8] (using early-layer attention), SparseVLM [60] (using text–to-vision cross-attention), DivPrune [2] (using diversity-driven selection), and HoloV [64] (using crop-wise attention and diversity). Another line of calibration-based approaches (e.g., FitPrune [57], VTW [31]) select pruning layers/ratios by analyzing model outputs on a calibration set. However, the first line relies mainly on semantic salience and overlooks the action-dependent requirements of VLA, leading to performance drops (see Sec. 3.2). The second line depends on offline calibration that is unfriendly to resource-constrained VLA deployment and still hinges on semantic salience. These limitations motivate a VLA-specific pruning method.
2.3 Training-Free Acceleration for VLA Models
@@
Visual token pruning has recently been integrated into training-free VLA acceleration frameworks [58]. For example, EfficientVLA [58], SP-VLA [29], and SpecPrune-VLA [50] perform structured acceleration that incorporates token pruning. However, these methods still use semantic salience to retain tokens. Thus, they rely on additional modules (e.g., diffusion-feature cache [56], lightweight generation module [29], action-aware controller [50], or layer-reduction [56, 50]) to balance performance and speed, limiting their generalizability and effectiveness. A recent method, VLA-Cache [55], exploits temporal continuity by caching static visual token features, opening a promising direction. However, its cache-based mechanism is less efficient than direct pruning and still depends on text-to-vision cross-attention to preserve task-relevant tokens in a coarse-grained manner. The above limitations stem from overlooking the dual-system nature of VLA models (see Sec. 3.2). In this paper, we show that token pruning can achieve significant VLA acceleration without sacrificing performance.
3 Preliminary Analysis
3.1 Background
@@
Vision-Language-Action (VLA) models process sequential inputs of the form into across timesteps . Here, represents the input visual observations, denotes language task instructions, and denotes the predicted action tokens, which can be detokenized into executable robotic actions. The text input is tokenized into text tokens . The vision encoder processes into image features, which are projected into visual tokens using a projector. These tokens, along with a proprioceptive token representing robot state, form the Transformer input. Since self-attention scales quadratically with token count, the (generally ) visual tokens dominate inference cost.
@@
(2) Action decode Stage. In this stage, queries are action tokens ( for autoregressive [22], for chunk-based [4]). Attention is computed as
@@
and quantifies each patch’s significance to semantic understanding [55, 60] and action execution, respectively.
3.2 Motivating Observation for VLA-Pruner
We systematically delve into the intrinsic characteristics of VLA inference (Fig. 2). Our findings identify the fundamental limitations of current VLM-specific token pruning methods for VLA models, which motivate our VLA-Pruner.
@@ 3.2 Motivating Observation for VLA-Pruner
As noted in previous works [60, 8], incorporation of visual tokens inevitably introduces substantial memory and computational overhead when compared to LLMs. For VLA models, the number of processed visual tokens (typically 256 per frame, where is the number of camera views) is often an order of magnitude larger than that of text tokens (30–50) and action tokens (7–56), making input visual stream prefilling the primary contributor to the computational cost. Since visual representations are highly redundant [2, 8, 60], visual token pruning shows great promise for VLA acceleration.
Motivation 2: Visual token pruning must account for the dual-system nature of VLA models.
The efficacy of VLA models stems from their inherent dual-system ability to integrate high-level semantic understanding and task planning with low-level action execution [13, 18, 19, 48, 11]. Our key insight is that this inherent duality manifests directly in VLA model inference, resulting in distinct visual requirements for its two stages: vision-language prefill and action decode. To investigate this, we analyzed the attention distributions of OpenVLA [22] on the LIBERO benchmark [32] (Fig. 2). The results reveal that VLA inference demonstrates distinct attention patterns: prefill attention exhibits a broad, semantic distribution, while action decode attention shows localized focus essential for precise motor control, as exemplified in Fig. LABEL:fig:prefill_attn and Fig. LABEL:fig:decode_attn. Quantitatively, the overlap ratio of top-attended visual tokens between vision–language prefilling and action decoding averages 50% and often falls below 30% (Fig. 2). Attention scores are averaged across all Transformer layers. Existing pruning strategies overlook this crucial attention divergence, thereby risking the premature removal of tokens vital for action generation but less salient during prefilling, particularly at high compression ratios. This oversight is critical, since removing even a few high-attention tokens is known to severely degrade performance [61, 16]. We also provide a qualitative illustration in Fig. 7.
@@
Since vision token prefilling dominates computational overhead (Motivation 1), token pruning must operate during early-layer prefilling [8]. However, while an ideal strategy should consider both semantic and action-level token importance (Motivation 2), action decode attention scores (practical proxy for action-level importance) are unavailable at prefill time. Fortunately, VLA robot manipulation exhibits temporal continuity [55]. We observe that action-to-vision attention maps show strong short-term temporal correlation: the top-attended visual tokens during action decoding at timesteps and show consistently high overlap ratios (Fig. 2 show ratios averaged across LIBERO tasks and a representative rollout). This temporal continuity makes statistical temporal smoothing techniques [52, 14, 21] an practical and effective proxy for estimating current action-to-vision attention.

### v3

@@ Abstract
Vision-Language-Action (VLA) models have shown promise for embodied AI but suffer from heavy computational overhead of processing continuous visual streams. Visual token pruning—keeping salient tokens and dropping redundant ones—emerges as a mainstream technique to accelerate Vision-Language Models (VLMs), offering a natural solution. However, these VLM-specific token pruning methods rely solely on semantic salience metrics (e.g., prefill attention) to select tokens, while overlooking VLA models’ intrinsic dual-system nature of high-level semantic understanding and low-level action execution. Consequently, they bias token retention toward semantic cues, discard critical information for action generation, and significantly degrade VLA performance. To bridge this gap, we propose VLA-Pruner, a plug-and-play VLA token pruning method that aligns with the dual-system nature of VLA models and exploits the temporal continuity in robot manipulation. Specifically, VLA-Pruner introduces a dual-objective importance criterion for token retention: vision-language prefill attention for semantic relevance and action decode attention, estimated via temporal smoothing, for execution precision. Based on this criterion, VLA-Pruner proposes a dual-level token selection strategy that adaptively preserves a compact set of visual tokens for both semantic understanding and action execution under given compute budget. Experiments show that VLA-Pruner outperforms state-of-the-art approaches across multiple VLA architectures, achieving up to 1.99 speedup with comparable manipulation quality. Our code is publicly available at: https://github.com/MINT-SJTU/VLA-Pruner.git.
@@ 1 Introduction
Recent advances in robot learning have led to remarkable progress in developing generalist policies capable of acting across diverse tasks and environments (Ma et al., 2024; Firoozi et al., 2025; Zhao et al., 2025; Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024; Black et al., 2024; Li et al., 2024a; Chi et al., 2025). Vision-Language-Action (VLA) models (Brohan et al., 2022; Zitkovich et al., 2023; Kim et al., 2024, 2025a; Black et al., 2024; Li et al., 2024a; Lin et al., 2025a), which benefit from the multimodal understanding and transfer ability of Vision-Language Models (VLMs) (Liu et al., 2023b; Bai et al., 2023), have emerged as a promising pathway toward general-purpose robot policy. Leveraging large-scale real-world robotic datasets (O’Neill et al., 2024), VLA models can perceive visual scenes, understand natural language instructions, and directly execute low-level actions to complete complex tasks across diverse scenarios. However, in real-time deployment, VLA models must process continuous visual streams within dynamic environments and thus struggle with heavy computational overhead. To mitigate this, existing works have explored acceleration techniques such as model lightweighting (Wen et al., 2025), quantization (Park et al., 2024), and early-exit (Yue et al., 2024) for VLA. Despite effectiveness to some extent, these methods require either architectural modifications or retraining, hindered by poor generalizability.
@@
Previous research (Chen et al., 2024; Shang et al., 2024; Alvar et al., 2025) has shown that visual representations (i.e., visual tokens) in Vision-Language Models (VLMs) are highly redundant. Therefore, visual token pruning, which retains the most salient tokens while pruning the rest, has become a general and effective approach to improve VLM inference efficiency (Alvar et al., 2025; Chen et al., 2024; Zhang et al., 2024; Shang et al., 2024; Ye et al., 2025; Xing et al., 2024; Zou et al., ) and shows promise for general VLA acceleration. These methods typically define importance criteria for tokens, such as attention scores (Chen et al., 2024; Zhang et al., 2024; Zou et al., ), gradient information (Mao et al., 2025a, b) or feature diversity (Alvar et al., 2025) to quantify the significance of visual tokens, and less important tokens are pruned at inference time. By reducing visual redundancy, they yield substantial compute savings with minimal accuracy loss on standard VLM tasks such as image captioning (Agrawal et al., 2019; Plummer et al., 2015), VQA (Hudson and Manning, 2019; Liu et al., 2024), and video understanding (Xu et al., 2017; Li et al., 2016).
@@
Despite recent progress, existing token pruning methods rely solely on semantic salience metrics measured during context prefilling—e.g., prefill attention in FastV (Chen et al., 2024) or text-to-vision cross-attention scores in SparseVLM (Zhang et al., 2024)—to rank visual tokens and prune the lowest-scoring ones. While such semantic-only criteria can be effective for VLMs, they are misaligned with the intrinsic dual-system nature of VLA models (Han et al., 2024; Huang et al., 2025; Sloman, 1996). Specifically, VLA inference couples high-level semantic understanding, task planning with low-level action execution, which entails distinct visual requirements between its vision-language prefill and action decode stages; correspondingly, we observe clearly different attention patterns across these two stages (see Sec. 3). Consequently, current methods tend to bias token retention towards semantic cues, thereby discarding critical details for action generation and significantly degrading VLA performance, especially at high pruning ratios (see Figure 1), as also observed in prior work (Xu et al., 2025). Thus, existing VLM-specific token pruning strategies are fundamentally ill-suited for VLA models, highlighting the urgent need for a VLA-specific method.
@@
To bridge this gap, we propose VLA-Pruner, a versatile plug-and-play token pruning approach that aligns with the dual-system nature of VLA models and exploits the temporal continuity in robot manipulation. VLA-Pruner offers two principal insights for VLA token pruning. First, VLA-Pruner adopts a dual-objective importance criterion for visual token retention: vision-language prefill attention scores for semantic relevance (following prior works (Chen et al., 2024; Ye et al., 2025; Mao et al., 2025a; Zhang et al., 2024)) and action decode attention scores for execution precise. A critical challenge to achieve this is that decode attention is unavailable at prefill time. Fortunately, we observe that temporal continuity of robot manipulation (Xu et al., 2025) manifests in VLA action attention patterns evolving: top-attended visual patches overlap substantially across consecutive timesteps (see Sec. 3). Leveraging this continuity, we practically estimate the current action-to-vision attention from recent decode attentions via statistical temporal smoothing techniques. Second, to ensure robust performance, our dual-level token selection strategy avoids fragile attention score fusion that incurs redundancy and bias. We draw on the minimal-redundancy-maximal-relevance (mRMR) principle (Peng et al., 2005) and adopt a patch-wise scheme: take the union of patches salient at either semantic or action level to maximize relevance, and then filter them by reducing redundancy, yielding a compact yet informative visual token set that meets the given compute budget. This innovative design reduces computational overhead and adaptively preserves essential information for both semantic understanding and action execution.
We summarize our contributions. (1) We reveal the dual-system property of VLA inference, observing distinct attention patterns between prefill and decode stages, which identifies the fundamental limitations of current VLM pruning methods for VLA acceleration (see Sec 3). (2) We propose VLA-Pruner, a general, training-free framework that aligns with this intrinsic duality and exploits the temporal continuity in VLA manipulation (see Sec 4). (3) Serving as a plug-and-play module for multiple VLA models, VLA-Pruner achieves state-of-the-art performance across two simulated environments, delivering up to 1.99 speedup with negligible performance drops and maintaining robustness even at an 87.5% pruning ratio (see Sec 5). We also valid its real-world advantage on a 6-DoF xArm6 robot.
2 Preliminary
2.1 Related Works
Vision-Language-Action Models (VLA)
@@ 2.1 Related Works
Large-scale vision–language models (VLMs) have propelled multimodal learning by incorporating visual perception and language reasoning (Li et al., 2023). Building on these advances, Vision–Language–Action (VLA) models (Kim et al., 2024; Black et al., 2024; Kim et al., 2025b; Brohan et al., 2022; Zitkovich et al., 2023) introduce an action modality, enabling end-to-end motor control. These models typically adopt large VLM backbones (Touvron et al., 2023) and are fine-tuned on robotic data (O’Neill et al., 2024). For action generation, implementations range from discretizing actions into language-like tokens (e.g., OpenVLA (Kim et al., 2024)) to attaching diffusion-policy heads (e.g., (Black et al., 2024)), typically relying on action-to-vision cross-attention for fine-grained interaction. Despite effectiveness on robot manipulation (Liu et al., 2023a; Li et al., 2024b), VLA models remain computationally intensive, hindering real-time deployment on resource-constrained platforms.
Visual Token Pruning for VLMs
@@
Visual token pruning (Chen et al., 2024; Zhang et al., 2024; Alvar et al., 2025; Shang et al., 2024; Ye et al., 2025) is a widely adopted strategy to reduce visual redundancy in VLMs. One line of these methods defines semantic importance criteria to rank and prune visual tokens, e.g., FastV (Chen et al., 2024) (using prefill attention), SparseVLM (Zhang et al., 2024) (using text–to-vision attention), DivPrune (Alvar et al., 2025) (using diversity-driven selection), and HoloV (Zou et al., ) (using crop-wise attention and diversity). Another line of calibration-based approaches (e.g., FitPrune (Ye et al., 2025), VTW (Lin et al., 2025b)) select pruning layers/ratios by analyzing model outputs on a calibration set. However, the first line relies mainly on semantic salience and overlooks the action-dependent requirements of VLA (see Sec. 3). The second line depends on offline calibration that is unfriendly to resource-constrained VLA deployment and still hinges on semantic salience. These limitations motivate a VLA-specific token pruning method.
Training-Free Acceleration for VLA Models
@@
Visual token pruning has recently been integrated into training-free VLA acceleration frameworks (Yu et al., 2025). To name a few, EfficientVLA (Yu et al., 2025), SP-VLA (Li et al., 2025), and SpecPrune-VLA (Wang et al., 2025) perform structured acceleration that incorporates token pruning. However, these methods still use semantic salience to retain tokens. Thus, they rely on additional modules (e.g., diffusion-feature cache (Yang et al., 2025), lightweight generation module (Li et al., 2025), or layer-reduction (Yang et al., 2025; Wang et al., 2025)) to balance performance and speed, limiting generalizability and effectiveness. A recent method, VLA-Cache (Xu et al., 2025), exploits temporal continuity by caching static visual token features, opening a promising direction. However, its cache-based mechanism is less efficient than direct pruning and still depends on text-to-vision cross-attention to preserve task-relevant tokens. The above limitations stem from overlooking the dual-system nature of VLA models (see Sec. 3). In this paper, we show that visual token pruning can achieve significant VLA inference acceleration without sacrificing performance.
2.2 Background of VLA Inference
@@
Vision-Language-Action (VLA) models process sequential inputs of the form into across timesteps . represents the input visual observations, denotes language task instructions, and denotes the predicted action tokens, which are detokenized into executable robotic actions. The text input is tokenized into text tokens . The vision encoder processes into image features, which are projected into visual tokens . These tokens, along with a proprioceptive token representing robot state, form the Transformer input. Since self-attention scales quadratically with token counts, the (generally ) visual tokens dominate inference cost.
@@
(2) Action decode Stage. In this stage, queries are action tokens ( for autoregressive (Kim et al., 2024), for chunk-based (Black et al., 2024)). Attention is computed as
@@
3 Motivation
We systematically delve into the intrinsic characteristics of VLA inference (Figure 2). Our findings identify the fundamental limitations of current VLM-specific token pruning methods for VLA models, which motivate our VLA-Pruner.
@@ 3 Motivation
As noted in previous works (Zhang et al., 2024; Chen et al., 2024), incorporation of visual tokens inevitably introduces substantial memory and computational overhead. For VLA models, the number of processed visual tokens (typically 256 per frame, where is the number of camera views) is often an order of magnitude larger than that of text tokens (30–50) and action tokens (7–56), making input visual stream prefilling the primary contributor to the computational cost. Since visual representations are highly redundant (Alvar et al., 2025; Chen et al., 2024; Zhang et al., 2024), visual token pruning shows great promise for VLA acceleration.
Motivation 2: Visual token pruning must account for the inherent dual-system nature of VLA models.
The efficacy of VLA models stems from their inherent dual-system ability to integrate high-level semantic understanding and task planning with low-level action execution (Firoozi et al., 2025; Han et al., 2024; Huang et al., 2025; Sloman, 1996; Cui et al., 2025). Our key insight is that this inherent duality manifests directly in VLA model inference, resulting in distinct visual requirements for its two stages: vision-language prefill and action decode. To investigate this, we analyzed the attention distributions of OpenVLA (Kim et al., 2024) on the LIBERO benchmark (Liu et al., 2023a) (Figure 2). The results reveal that VLA inference demonstrates distinct attention patterns: prefill attention exhibits a broad, semantic distribution, while action decode attention shows localized focus essential for precise motor control, as exemplified in Figure LABEL:fig:prefill_attn and Figure LABEL:fig:decode_attn. Quantitatively, the overlap ratio of top-attended visual tokens between vision-language prefilling and action decoding averages 50% and often falls below 30% (Figure 2). Attention scores are averaged across all Transformer layers. Existing pruning strategies overlook this crucial attention divergence, thereby risking the premature removal of tokens vital for action generation but less salient during prefilling, particularly at high compression ratios. This oversight is critical, since removing even a few high-attention tokens is known to severely degrade performance (Zhang et al., 2023; Ge et al., 2023). We provide a qualitative illustration in Figure 7.
@@
Since vision token prefilling dominates computational overhead (Motivation 1), token pruning must operate during early-layer prefilling (Chen et al., 2024). However, while an ideal strategy should consider both semantic and action-level token importance (Motivation 2), action decode attention scores (practical proxy for action-level importance) are unavailable at prefill time. Fortunately, VLA robot manipulation exhibits temporal continuity (Xu et al., 2025). We observe that action-to-vision attention maps show strong short-term temporal correlation: top-attended visual tokens during action decoding at timesteps and show consistently high overlap ratios. Figure 2 shows averaged ratios and a representative rollout. This continuity makes statistical temporal smoothing techniques (Winters, 1960; Gardner, 1985; Hyndman and Athanasopoulos, 2018) a practical and effective proxy for estimating current action-to-vision attention.